OpenStudy (danysanchy97):

Divide 6/13 by 6/12 . A. 1/12 B. 13/12 C. 9/16 D. 12/13

OpenStudy (danysanchy97):

is it D?

OpenStudy (anitas.):

\[\frac{ 6 }{ 13 }/\frac{ 6 }{ 12 }=\frac{ 6 }{ 13}*\frac{ 12 }{ 6}=\frac{ 72 }{ 78 }=\frac{ 12 }{ 13}\] Therefore your answer is D. 12/13

OpenStudy (danysanchy97):

thank you can you help me on some more?

OpenStudy (anitas.):

sure open new posts and tag me ok?

jhonyy9 (jhonyy9):

@AnitaS. sorry but without multiplie them you can simplifie 6 from numerator by 6 from denominator and will get the right answer - yes ?